OpenStudy (anonymous):

how do we write this in standard form 1.2345 x 10 ^8 ?

OpenStudy (anonymous):

do they show u how to do it ?

OpenStudy (anonymous):

do you know what number is 10^8?

OpenStudy (anonymous):

or let's say what number is 10^2?

OpenStudy (anonymous):

1.2345 x 10 ^8 To get the standard form, just expand the 10^8 and then multiply 10^8 = 10 x 10 x 10 x 10 x 10 x 10 x 10 x 10 = 100000000 so 1.2345 x 10 ^8 = 1.2345 x 100000000 = 123450000.0000 now just ignore the decimal and the zeros after that and you get the final answer 1.2345 x 10 ^8 = 123450000

OpenStudy (anonymous):

it's gona b 123,450,000 or something. all u have 2 do is take your pencil and do this :  excuse the handwriting :)